# Higher-Order Corrections to Scrambling Dynamics in Brownian Spin SYK Models

### v1

@@ 1 Introduction
Scrambling refers to the dynamical process in quantum many-body systems whereby initially localized information becomes highly nonlocal, encoded in increasingly complex correlations. Consequently, this information becomes effectively inaccessible to local measurements despite being preserved under unitary time evolution. A central quantitative probe of scrambling Lewis-Swan:2019xbi ; Xu_2024 is provided by out-of-time-order correlators (OTOCs) Larkin1969QuasiclassicalMI , which diagnose the growth of noncommutativity between initially simple operators. Microscopically, this process is described as operator growth: under Heisenberg picture, a simple operator spreads into a superposition of increasingly nonlocal operators Nahum_2018 ; vonKeyserlingk2018 ; Parker_2019 , reflecting the delocalization of quantum information across the system. While operator growth is dual to entanglement generation in the Schrödinger picture Kim:2013etb ; Ho_2017 ; Nahum_2017 ; swann2023spacetimepictureentanglementgeneration , the operator-size distribution offers a more fine-grained probe of scrambling dynamics than the mean size alone. It captures the full spectrum of operator growth during unitary time evolution Roberts_2018 ; Qi_2019 ; Lucas_2020 ; Zhang_2023 as well as in open quantum systems Zhangoperator_2023 ; Schuster2023 , and has also been addressed in experimental studies qi2019measuringoperatorsizegrowth .
In realistic experimental settings, however, probing scrambling is complicated by the presence of noise and experimental imperfections. While unitary evolution delocalizes information via operator growth, decoherence and imperfect time reversal can obscure or even mimic the decay of OTOCs expected from intrinsic unitary dynamics. Motivated by this challenge, Nadie2025 introduced a protocol based on dressed OTOCs and an echo observable, which explicitly incorporates unequal forward and backward evolutions as well as depolarizing noise. By studying the ratio between the dressed OTOC and the echo signal (termed the renormalized OTOC, or ROTOC, and defined in PhysRevA.97.062113 ), that work provided a framework for disentangling genuine scrambling dynamics from experimental imperfections.
@@
Here $\mathbb{E}$ denotes the ensemble average. The Brownian spin SYK model is a type of quantum many-body system known as Brownian models Lashkari2013 ; saad2019semiclassicalrampsykgravity ; Jian_2021 ; Stanford_2022 , characterized by its inclusion of time-dependent, stochastic couplings that act as a source of dephasing noise. A Similar model is also discussed in Erd_s_2014 ; Baldwin_2020 ; Berkooz_2018 ; Sunderhauf:2019djv ; Yin:2020pjd ; Swingle:2023nvv ; Hanada_2024 ; anschuetz2024boundsgroundstateenergy ; Xu_2025 as spin SYK model which served as an extension of the conventional SYK model made by fermions Kitaev2015 ; Polchinski_2016 ; Maldacena_2016 ; Jevicki:2016bwu ; jevicki2016bilocalholographysykmodel .
@@
Our approach is formulated in the large-$N$ limit. We first extend the dimension of the original matrix $M$ formally to infinity, obtaining $M_{\infty}$, whose entries retain their explicit dependence on the system parameter $N$. We then define an approximate generating function $\mathsf{G}(x,t)$ by
@@ 2.1 Brownian spin SYK model and operator dynamics
$O_{A_{n}}=\bigotimes_{k=1}^{n}\sigma_{j_{k}},\sigma_{j}\in\{X,Y,Z\}\,.$ (11)
@@
$\mathbb{E}(J_{A_{n}}(t)J_{A_{n}^{\prime}}(t^{\prime}))=\delta_{A_{n}A_{n}^{\prime}}\delta(t-t^{\prime})\mu_{n}^{2}\,.$ (12)
@@
Under unitary evolution, this distribution is normalized, $\sum_{w}b^{\text{clean}}_{w}(t)=1$, reflecting the preservation of the operator’s trace.
2.2 Construct ROTOC with $b_{w}(t)$
@@
$\mathbb{E}(\widetilde{J}_{A}(t)\widetilde{J}_{B}(t^{\prime}))=r\delta_{AB}\delta(t-t^{\prime})\mu_{\mathsf{w}(O_{A})}^{2}\,.$
@@ 3.2 Leading order solution
Once the generating function $\mathsf{G}^{(0)}(x,t)$ is determined, physical observables follow directly. For example, $\langle w\rangle_{c}$ used in the calculation of the OTOC can be readily obtained as

### v2

@@ 1 Introduction
Scrambling refers to the dynamical process in quantum many-body systems whereby initially localized information becomes highly nonlocal, encoded in increasingly complex correlations. Consequently, this information becomes effectively inaccessible to local measurements despite being preserved under unitary time evolution. A central quantitative probe of scrambling Lewis-Swan:2019xbi ; Xu_2024 is provided by out-of-time-order correlators (OTOCs) Larkin1969QuasiclassicalMI , which diagnose the growth of noncommutativity between initially simple operators. Beyond their foundational role, OTOCs have been linked to entanglement entropy FAN2017707 and have been experimentally measured on various platforms Li_2017 ; PhysRevLett.120.070501 ; PhysRevA.100.013623 ; Gu_2022 ; Li_2024 . Microscopically, this process is described as operator growth: in the Heisenberg picture, a simple operator spreads into a superposition of increasingly nonlocal operators Nahum_2018 ; vonKeyserlingk2018 ; Parker_2019 , reflecting the delocalization of quantum information across the system. While operator growth is dual to entanglement generation in the Schrödinger picture Kim:2013etb ; Ho_2017 ; Nahum_2017 ; swann2023 , the operator-size distribution offers a more fine-grained probe of scrambling dynamics than the mean size alone. It captures the full spectrum of operator growth during unitary time evolution Roberts_2018 ; Qi_2019 ; Lucas_2020 ; Zhang_2023 as well as in open quantum systems Zhangoperator_2023 ; Schuster2023 , and has also been addressed in experimental studies qi2019measuringoperatorsizegrowth .
In realistic experimental settings, however, probing scrambling is complicated by the presence of noise and experimental imperfections. While unitary evolution delocalizes information via operator growth, decoherence and imperfect time reversal can obscure or even mimic the decay of OTOCs expected from intrinsic unitary dynamics. Motivated by this challenge, Nadie2025 introduced a protocol based on dressed OTOCs and an echo observable, which explicitly incorporates unequal forward and backward evolutions as well as depolarizing noise. [^1] The influence of noise on OTOCs and quantum chaos diagnostics has also been examined in Li_2026 and Li:2025gil , respectively. By studying the ratio between the dressed OTOC and the echo signal (termed the renormalized OTOC, or ROTOC, and defined in PhysRevA.97.062113 ), that work provided a framework for disentangling genuine scrambling dynamics from experimental imperfections.
@@
Here $\mathbb{E}$ denotes the ensemble average. The Brownian spin SYK model is a type of quantum many-body system known as Brownian models Lashkari2013 ; saad2019semiclassicalrampsykgravity ; Jian_2021 ; Stanford_2022 , characterized by its inclusion of time-dependent, stochastic couplings that act as a source of dephasing noise. A similar model is also discussed in Erd_s_2014 ; Baldwin_2020 ; Berkooz_2018 ; Sunderhauf:2019djv ; Yin:2020pjd ; Swingle:2023nvv ; Hanada_2024 ; anschuetz2024boundsgroundstateenergy ; Xu_2025 ; basu2025complexityquadraticquantumchaos as spin SYK model which served as an extension of the conventional SYK model made by fermions Kitaev2015 ; Polchinski_2016 ; Maldacena_2016 ; Jevicki:2016bwu ; jevicki2016bilocalholographysykmodel .
@@
Our approach is formulated in the large-$N$ limit. We first extend the dimension of the original matrix $M$ formally to infinity, obtaining $M_{\infty}$, whose entries retain their explicit dependence on the system parameter $N$. We then define an approximate generating function $\mathsf{G}(x,t)$ by  [^3] In Appendix of Nadie2025 , the author employs the generating function technique to derive the leading-order results. Meanwhile, in Zhang:2022fma , the authors first consider the large $N$ limit and subsequently take the continuous limit, after which the generating function method is introduced.
@@ 2.1 Brownian spin SYK model and operator dynamics
$O_{A_{n}}=\bigotimes_{k=1}^{n}\sigma_{j_{k}},\sigma_{j_{k}}\in\{X,Y,Z\}\,.$ (11)
@@
$\mathbb{E}(J_{A_{n}}(t)J_{A_{m}^{\prime}}(t^{\prime}))=\delta_{A_{n}A_{m}^{\prime}}\delta(t-t^{\prime})\mu_{n}^{2}\,.$ (12)
@@
Under unitary evolution, this distribution is normalized $\sum_{w}b^{\text{clean}}_{w}(t)=1$, reflecting the preservation of the operator’s trace.
2.2 Construct ROTOC with operator size distribution $b_{w}$
@@
$\mathbb{E}(J_{A}(t)\widetilde{J}_{B}(t^{\prime}))=r\delta_{AB}\delta(t-t^{\prime})\mu_{\mathsf{w}(O_{A})}^{2}\,.$
@@ 3.2 Leading order solution
Once the generating function $\mathsf{G}^{(0)}(x,t)$ is determined, physical observables follow directly. For example, $\langle w\rangle_{c}$ used in the calculation of OTOC can be easily obtained as